### 25 Apr 2026 v1

# A proof of Bala’s general-$m$ representation of the harmonic numbers

Abstract.
For every nonzero integer $m$ and every integer $n\geq 1$, the $n$th harmonic number $H_{n}=1+\tfrac{1}{2}+\dots+\tfrac{1}{n}$ satisfies the identity
$H_{n}\;=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
The cases $m=1$ and $m=2$ are classical; for general nonzero integer $m$ the identity was conjectured by P. Bala in the OEIS entry A001008 in 2022 and remained open. We prove it here, working throughout in $\mathbb{Q}[[x]]$. The proof reduces, via a substitution $u=x/(1-x)^{m}$, to two formal-power-series identities: a Lagrange–Bürmann evaluation of $\sum_{k\geq 1}\binom{mk}{k}u^{k}/k$, and the fixed-point fact that under that substitution the unique solution $v(u)$ of $v=u(1-v)^{m}$ is $v=x$. The argument extends verbatim to arbitrary complex $m\neq 0$.
Key words and phrases:
Harmonic number; binomial sum; Lagrange–Bürmann formula; formal power series; Fuss–Catalan; OEIS A001008
2020 Mathematics Subject Classification:
05A19, 05A10, 11B65, 11B83
1. Introduction
The numerators of the harmonic numbers
$H_{n}\;:=\;\sum_{i=1}^{n}\frac{1}{i},\qquad n\geq 1,\quad H_{0}:=0,$
form the integer sequence OEIS A001008. The On-Line Encyclopedia of Integer Sequences (OEIS) hosts a number of binomial-coefficient representations of $H_{n}$ contributed over time as comments to that entry. One of these, contributed by P. Bala on March 4, 2022, is the following parametric family.
Theorem 1 (Bala’s identity, conjectured 2022).
For every nonzero integer $m$ and every integer $n\geq 1$,
(1) $H_{n}\;=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
[…]
$H_{n}=\tfrac{1}{2}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{2k}{k}\binom{n+k}{n-k}$,
was contributed (also as a conjecture) to A001008 by G. Detlefs in April 2013 and proved shortly thereafter. For $m\geq 3$ and for every negative integer $m$, the identity (1) appears to have remained open until now.
In this note we prove Theorem 1 and observe that the proof in fact establishes the identity for every complex $m\neq 0$ when both sides are read as polynomials in $m$ for fixed $n$ (Remark 5).
The proof is via formal power series. In Section 2 we sum the right-hand side of (1) as an ordinary generating function $A_{m}(x)=\sum_{n}S_{m}(n)\,x^{n}$ and reduce it, under the substitution $u=x/(1-x)^{m}$, to a sum $T(u)=\sum_{k\geq 1}(-1)^{k+1}\binom{mk}{k}u^{k}/k$. Section 3 evaluates $T(u)$ via a Lagrange–Bürmann calculation. Section 4 closes the loop with a formal fixed-point lemma: under $u=x/(1-x)^{m}$, the unique solution $v(u)$ of $v=u(1-v)^{m}$ in $x\,\mathbb{Q}[[x]]$ is $v=x$. The conclusion $A_{m}(x)=-\log(1-x)/(1-x)=\sum_{n}H_{n}\,x^{n}$ is then immediate.
Throughout, we use the generalized binomial coefficient $\binom{a}{b}=a(a-1)\cdots(a-b+1)/b!$ for $a\in\mathbb{Z}$, $b\in\mathbb{Z}_{\geq 0}$, so that (1) is meaningful for all nonzero integer $m$. For negative $m$ this binomial may take negative or zero values term by term, but the sum still equals $H_{n}$ (see Example 7).
2. The generating function and a substitution
For nonzero integer $m$, write $S_{m}(n)$ for the right-hand side of (1), with the convention $S_{m}(0)=0$:
$S_{m}(n)\;:=\;\frac{1}{m}\,\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{mk}{k}\binom{n+(m-1)k}{n-k}$.
Let
$A_{m}(x)\;:=\;\sum_{n\geq 0}S_{m}(n)\,x^{n}\;\in\;\mathbb{Q}[[x]].$
Theorem 1 is equivalent to the formal power-series identity $A_{m}(x)=-\log(1-x)/(1-x)$, since $\log(1-x)/(1-x)=\sum_{n\geq 1}H_{n}\,x^{n}.$ We compute $A_{m}(x)$ in two steps.
Step 1: switch order of summation. We have
(2) $A_{m}(x)$ $=\frac{1}{m}\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\sum_{n\geq k}\binom{n+(m-1)k}{n-k}\,x^{n}.$
Setting j=nkj=n-k in the inner sum and using $\sum_{j\geq 0}\binom{j+r}{j}\,x^{j}=(1-x)^{-r-1}$ with $r=mk$,
(3) $\sum_{n\geq k}\binom{n+(m-1)k}{n-k}\,x^{n}=x^{k}\sum_{j\geq 0}\binom{j+mk}{j}\,x^{j}=\frac{x^{k}}{(1-x)^{mk+1}}.$
Step 2: introduce $u=x/(1-x)^{m}$. Substituting (3) into (2) and pulling out the factor $(1-x)^{-1}$,
(4) $A_{m}(x)=\frac{1}{m\,(1-x)}\,\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\,u^{k},\qquad u\;:=\;\frac{x}{(1-x)^{m}}.$
The proof of Theorem 1 therefore reduces to evaluating the inner sum in (4) as a formal power series in $u$, and then identifying the result after $u\mapsto x/(1-x)^{m}$.
3. A Lagrange–Bürmann identity
The next lemma is the key formal-series identity. Its proof uses the Lagrange–Bürmann formula in the coefficient-extraction form (see e.g. Stanley [3, Theorem 5.4.2]).
Lemma 2.
For every nonzero integer $m$,
(5) $\sum_{k\geq 1}\frac{1}{k}\binom{mk}{k}\,u^{k}\;=\;m\,\log(1+w),$
where $w=w(u)\in u\,\mathbb{Q}[[u]]$ is the unique formal power series satisfying
(6) $w\;=\;u\,(1+w)^{m}.$
Equivalently,
(7) $\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\binom{mk}{k}\,u^{k}\;=\;-\,m\,\log(1-v),$
where $v=v(u)\in u\,\mathbb{Q}[[u]]$ is the unique formal power series satisfying $v=u\,(1-v)^{m}$.
Proof.
The implication (5) $\Rightarrow$ (7) is obtained by replacing $u$ by u-u and setting v=w(u)v=-w(-u), which maps (6) to $v=u(1-v)^{m}$ and $\log(1+w(-u))$ to $\log(1-v)$.
So we prove (5). Equation (6) has a unique solution $w\in u\,\mathbb{Q}[[u]]$ because $\phi(w):=(1+w)^{m}$ has $\phi(0)=1\neq 0$ — even when $m$ is a negative integer, in which case $(1+w)^{m}$ is read as the formal series $\sum_{j\geq 0}\binom{-|m|}{j}\,w^{j}$. This is the standard Lagrange-inversion hypothesis.
Differentiate (6) implicitly with respect to $u$:
$w^{\prime}=(1+w)^{m}+u\,m\,(1+w)^{m-1}\,w^{\prime}.$
Solving for $w^{\prime}$ and using $u\,(1+w)^{m-1}=w/(1+w)$ (which follows from (6) after dividing both sides by $1+w$),
(8) $w^{\prime}\;=\;\frac{(1+w)^{m}}{1-mw/(1+w)}\;=\;\frac{(1+w)^{m+1}}{1+(1-m)\,w}.$
By the Lagrange–Bürmann coefficient-extraction formula ([3, Theorem 5.4.2]), for $\phi(w)=(1+w)^{m}$ and any formal series $f$,
$\sum_{k\geq 0}u^{k}\,[w^{k}]\bigl(f(w)\,\phi(w)^{k}\bigr)\;=\;\frac{f(w)}{1-u\,\phi^{\prime}(w)}.$
Applying this with $f(w)=1$ and using $w^{k}]\bigl((1+w)^{mk}\bigr)=\binom{mk}{k}$,
(9) $\sum_{k\geq 0}\binom{mk}{k}\,u^{k}\;=\;\frac{1}{1-u\,m\,(1+w)^{m-1}}\;=\;\frac{1}{1-mw/(1+w)}\;=\;\frac{1+w}{1+(1-m)\,w},$
where the second equality is again $u(1+w)^{m-1}=w/(1+w)$.
Subtracting the $k=0$ term and dividing by $u$,
(10) $\sum_{k\geq 1}\binom{mk}{k}\,u^{k-1}\;=\;\frac{1}{u}\!\left(\frac{1+w}{1+(1-m)w}-1\right)\;=\;\frac{1}{u}\cdot\frac{m\,w}{1+(1-m)\,w}\;=\;\frac{m\,(1+w)^{m}}{1+(1-m)\,w},$
where the last equality uses $1/u=(1+w)^{m}/w$ from (6).
Set
$T(u)\;:=\;\sum_{k\geq 1}\frac{1}{k}\binom{mk}{k}\,u^{k}\;\in\;u\,\mathbb{Q}[[u]].$
By (10), $dT/du$ equals the right-hand side of (10). Combining with (8) and using the chain rule with $w=w(u)$,
$\frac{dT}{dw}\;=\;\frac{dT}{du}\cdot\frac{du}{dw}\;=\;\frac{m\,(1+w)^{m}}{1+(1-m)\,w}\cdot\frac{1+(1-m)\,w}{(1+w)^{m+1}}\;=\;\frac{m}{1+w}.$
Integrating with $T(0)=0$ and $w(0)=0$ gives $T(u)=m\,\log(1+w)$, which is (5). ∎
Remark 3.
For each fixed $m$ the series $w$ is a Fuss–Catalan-type series: $1+w$ is the generating function of $m$-ary trees (shifted), and (1+w)/(1+(1m)w)(1+w)/(1+(1-m)w) in (9) is the well-known generating function for the binomials $\binom{mk}{k}$. For $m=2$, (6) solves to $1+w=2/(1+\sqrt{1-4u})$ and Lemma 2 specialises to the classical identity $\sum_{k\geq 1}\binom{2k}{k}u^{k}/k=2\log\bigl(2/(1+\sqrt{1-4u})\bigr)$.
4. The fixed-point substitution
Lemma 2 evaluates the inner sum in (4) once $v(u)$ is computed under $u=x/(1-x)^{m}$. The next lemma is the identification we need.
Lemma 4.
Let $v(u)\in u\,\mathbb{Q}[[u]]$ be the unique formal solution of $v=u(1-v)^{m}$. Under the formal substitution $u=x/(1-x)^{m}$, one has v(u)=xv(u)\;=\;x as formal power series in $x$.
Proof.
Let $V(x):=v\!\bigl(x/(1-x)^{m}\bigr)\in x\,\mathbb{Q}[[x]]$ be the formal-series composition. Substituting $u=x/(1-x)^{m}$ into $v=u(1-v)^{m}$ shows that $V$ satisfies
(11) $V(x)\;=\;\frac{x}{(1-x)^{m}}\,\bigl(1-V(x)\bigr)^{m}.$
Define the map $G:x\,\mathbb{Q}[[x]]\to x\,\mathbb{Q}[[x]]$ by
$G(\xi)\;:=\;\frac{x}{(1-x)^{m}}\,(1-\xi)^{m},$
where $(1-\xi)^{m}$ is the binomial formal series $\sum_{j\geq 0}\binom{m}{j}(-\xi)^{j}$ when $m\geq 0$, and the analogous expansion $\sum_{j\geq 0}\binom{-|m|+j-1}{j}(-1)^{j}\xi^{j}$ when $m<0$. In either case $(1-\xi)^{m}\in 1+\xi\,\mathbb{Q}[[\xi]]$ for $\xi(0)=0$. So $G$ is well-defined.
$G$ is a strict contraction in the $x$-adic topology. Indeed, if $\xi_{1}\equiv\xi_{2}\pmod{x^{N}}$ for some $N\geq 1$, then $(1-\xi_{1})^{m}-(1-\xi_{2})^{m}\in x^{N}\,\mathbb{Q}[[x]]$ by direct binomial expansion (each coefficient of the difference is a polynomial in $\xi_{1}-\xi_{2}$ with no constant term). Multiplying by $x/(1-x)^{m}\in x\,\mathbb{Q}[[x]]$ shifts the order by one further, so
$G(\xi_{1})\;\equiv\;G(\xi_{2})\pmod{x^{N+1}}$.
By the formal Banach fixed-point theorem (the $x$-adic completion of $\mathbb{Q}[x]$ is complete), $G$ has a unique fixed point in $x\,\mathbb{Q}[[x]]$.
Exhibiting the fixed point. Take $\xi=x$. Then $G(x)=x\,(1-x)^{m}/(1-x)^{m}=x$, so $x$ is a fixed point of $G$. By uniqueness, $V(x)=x$. ∎
5. Proof of Theorem 1
Proof of Theorem 1.
Substitute (7) into (4) with the $v$ in (7) viewed under $u=x/(1-x)^{m}$. By Lemma 4 this $v$ equals $x$, hence $\log(1-v)=\log(1-x)$ and
$A_{m}(x)\;=\;\frac{1}{m\,(1-x)}\cdot\bigl(-m\,\log(1-x)\bigr)\;=\;\frac{-\log(1-x)}{1-x}.$
The right-hand side is the standard generating function $\sum_{n\geq 1}H_{n}\,x^{n}$, obtained by Cauchy-multiplying the two series $\log(1-x)=\sum_{n\geq 1}x^{n}/n$ and $1/(1-x)=\sum_{n\geq 0}x^{n}$. Comparing coefficients of $x^{n}$ on both sides of $A_{m}(x)=-\log(1-x)/(1-x)$ yields $S_{m}(n)=H_{n}$ for every $n\geq 1$, which is (1). ∎
6. Remarks and examples
Remark 5 (Extension to complex $m$).
For each fixed $n\geq 1$, both sides of (1) are rational functions of $m$ (after clearing the leading $1/m$, they are polynomials in $m$ of degree at most n1n-1). The proof above establishes equality of these rational functions on the infinite set $\mathbb{Z}\setminus\{0\}$ of nonzero integers. Two rational functions agreeing on an infinite set are equal as rational functions, so (1) holds for every complex $m\neq 0$ when both sides are interpreted as rational functions of $m$.
Example 6 (Small $m$).
The cases $m=1,2,3,4$ of (1) read
$m=1:\quad$ $H_{n}=\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{n}{k}$,
$m=2:\quad$ $H_{n}=\tfrac{1}{2}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{2k}{k}\binom{n+k}{n-k}$,
$m=3:\quad$ $H_{n}=\tfrac{1}{3}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{3k}{k}\binom{n+2k}{n-k}$,
$m=4:\quad$ $H_{n}=\tfrac{1}{4}\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\binom{4k}{k}\binom{n+3k}{n-k}$.
The first is the classical formula [1, (5.41)]; the second is Detlefs [6]; the rest are special cases of Theorem 1.
Example 7 (Negative $m$).
For m=1m=-1, the formula (1) becomes
$H_{n}\;=\;-\sum_{k=1}^{n}\frac{(-1)^{k+1}}{k}\,\binom{-k}{k}\binom{n-2k}{n-k}$,
where the binomials are the generalized binomials defined in Section 1. Most terms vanish because $\binom{n-2k}{n-k}=0$ for $\lceil n/2\rceil\leq k\leq n-1$ (the top index lies in $\{0,1,\dots,n-k-1\}$). For $n=4$ the only nonzero terms are $k=3$ and $k=4$, contributing $20/3$ and 35/4-35/4 respectively, with sum 25/12-25/12, so $(-25/12)=25/12=H_{4}$. The identity (1) has been numerically verified as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,\dots,8,10,15\}$ and $n\in\{1,\dots,40\}$ — all 640640 instances agree.
Remark 8 (Combinatorial interpretation).
The factor $\binom{mk}{k}/k$ has, up to a Fuss–Catalan factor of (m1)k+1(m-1)k+1, the meaning of the number of $m$-ary trees with $k$ internal nodes; cf. Stanley [3, §6.2] or Knuth [2, §7.2.1.6]. The factor $\binom{n+(m-1)k}{n-k}$ counts non-decreasing integer sequences of length nkn-k with values in $\{0,1,\dots,(m-1)k\}$. So (1) expresses $H_{n}$ as an inclusion–exclusion sum over a parameter $k$ that selects an $m$-ary tree size and a corresponding ”stretching” sequence. A purely bijective proof of (1) that does not pass through generating functions would be of independent interest.
Remark 9 (An infinite family of representations).
As $m$ ranges over $\mathbb{Z}\setminus\{0\}$, (1) gives an infinite family of distinct binomial-coefficient representations of $H_{n}$. By Remark 5, the family extends to a single algebraic relation in the parameter $m$. The ”right” coefficient identities — those obtained by extracting $m^{j}]$ from the polynomial identity in Remark 5 — encode binomial sums for the elementary symmetric polynomials in $\{1/1,1/2,\dots,1/n\}$ and may merit further investigation.
7. Numerical verification
The identity (1) and the intermediate identities (5), (7), and Lemma 4 have been verified by direct computation:
  • Both sides of (1) computed as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,3,4,5,6,7,8,10,15\}$ and $n\in\{1,2,\dots,40\}$ — all 640640 instances agree exactly.
  • The key Lagrange identity (5), its sign-alternating version (7), the classical evaluation $\sum_{k\geq 0}\binom{mk}{k}u^{k}=(1+w)/(1+(1-m)w)$, and the substitution $V(x)=x$ of Lemma 4 have all been checked as formal power-series identities to degree 88 in $u$ (resp. $x$) for $m\in\{-3,-2,-1,1,2,3,4,5\}$.
The verification scripts (Python; fractions.Fraction and sympy only) are available as supplementary material.
Acknowledgements
The author thanks Peter Bala for posing the conjecture proved here as a comment to OEIS sequence A001008, and the maintainers of the OEIS for hosting an indispensable repository of conjectured identities.

### 30 Apr 2026 v2

@@
 3. A Lagrange–Bürmann identity
-The next lemma is the key formal-series identity. Its proof uses the Lagrange–Bürmann formula in the coefficient-extraction form (see e.g. Stanley [3, Theorem 5.4.2]).
+The next lemma is the key formal-series identity.
+Its proof uses the Lagrange–Bürmann formula in the coefficient-extraction form (see e.g.
+Stanley [3, Theorem 5.4.2]; for a short elementary inductive proof of the Lagrange inversion formula, see Surya and Warnke [4]).
 Lemma 2.
@@
 For $n=4$ the only nonzero terms are $k=3$ and $k=4$, contributing $20/3$ and 35/4-35/4 respectively, with sum 25/12-25/12, so $(-25/12)=25/12=H_{4}$.
-The identity (1) has been numerically verified as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,\dots,8,10,15\}$ and $n\in\{1,\dots,40\}$ — all 640640 instances agree.
+The identity (1) has been numerically verified as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,\dots,8,10,15\}$ and $n\in\{1,\dots,40\}$ — all $640$ instances agree.
 Remark 8 (Combinatorial interpretation).
@@
 The factor $\binom{n+(m-1)k}{n-k}$ counts non-decreasing integer sequences of length nkn-k with values in $\{0,1,\dots,(m-1)k\}$.
-So (1) expresses $H_{n}$ as an inclusion–exclusion sum over a parameter $k$ that selects an $m$-ary tree size and a corresponding ”stretching” sequence.
+So (1) expresses $H_{n}$ as an inclusion–exclusion sum over a parameter $k$ that selects an $m$-ary tree size and a corresponding "stretching" sequence.
 A purely bijective proof of (1) that does not pass through generating functions would be of independent interest.
@@
 By Remark 5, the family extends to a single algebraic relation in the parameter $m$.
-The ”right” coefficient identities — those obtained by extracting $m^{j}]$ from the polynomial identity in Remark 5 — encode binomial sums for the elementary symmetric polynomials in $\{1/1,1/2,\dots,1/n\}$ and may merit further investigation.
+The "right" coefficient identities — those obtained by extracting $m^{j}]$ from the polynomial identity in Remark 5 — encode binomial sums for the elementary symmetric polynomials in $\{1/1,1/2,\dots,1/n\}$ and may merit further investigation.
 7. Numerical verification
-The identity (1) and the intermediate identities (5), (7), and Lemma 4 have been verified by direct computation:
-  • Both sides of (1) computed as exact rationals for $m\in\{-7,-5,-4,-3,-2,-1,1,2,3,4,5,6,7,8,10,15\}$ and $n\in\{1,2,\dots,40\}$ — all 640640 instances agree exactly.
-  • The key Lagrange identity (5), its sign-alternating version (7), the classical evaluation $\sum_{k\geq 0}\binom{mk}{k}u^{k}=(1+w)/(1+(1-m)w)$, and the substitution $V(x)=x$ of Lemma 4 have all been checked as formal power-series identities to degree 88 in $u$ (resp.
-$x$) for $m\in\{-3,-2,-1,1,2,3,4,5\}$.
-The verification scripts (Python; fractions.Fraction and sympy only) are available as supplementary material.
+The identity (1) and the four intermediate identities used in the proof — (5), (7), the classical evaluation (9), and the substitution $V(x)=x$ of Lemma 4 — have been verified by direct computation:
+  • Both sides of (1) computed as exact rationals (Python fractions.Fraction) for $m\in\{-7,-5,-4,-3,-2,-1,1,2,3,4,5,6,7,8,10,15\}$ and $n\in\{1,2,\dots,40\}$ — all $640$ instances agree exactly.
+See Appendix A.
+  • The four intermediate formal-power-series identities have all been checked coefficient-by-coefficient up to degree $8$ in $u$ (resp. $x$) for $m\in\{-3,-2,-1,1,2,3,4,5\}$. See Appendix B.
+The two verification scripts are self-contained Python (standard library only, no third-party dependencies) and reproduce in seconds. They are provided in full in Appendices A and B.
 Acknowledgements
 The author thanks Peter Bala for posing the conjecture proved here as a comment to OEIS sequence A001008, and the maintainers of the OEIS for hosting an indispensable repository of conjectured identities.
+The author is also grateful to Lutz Warnke for kindly drawing attention to the recent elementary inductive proof of the Lagrange inversion formula in [4].
+Appendix A Exact-rational verification of identity (1)
+The script below evaluates both sides of identity (1) as exact rationals (fractions.Fraction), without any floating point, for every pair $(m,n)$ with $m\in\{-7,-5,-4,-3,-2,-1,1,2,3,4,5,6,7,8,10,15\}$ and $n\in\{1,\dots,40\}$640640 instances.
+All instances match.
+Listing 1: verify_bala.py: exact-rational verification of (1).
+1"""
+2Verify Peter Bala's conjecture (2022) on A001008:
+3For any nonzero integer m,
+[…]
+5
+6where H(n) = 1 + 1/2 + ... + 1/n.
+7
+8Strategy: compute both sides as exact Fraction, compare equality.
+9"""
+10
+11from fractions import Fraction
+12from math import comb
+13
+14
+15def harmonic(n: int) -> Fraction:
+[…]
+17 h = Fraction(0)
+18 for i in range(1, n + 1):
+19 h += Fraction(1, i)
+20 return h
+21
+22
+23def gen_binom(top: int, k: int) -> int:
+24 """Compute generalized binomial coefficient C(top, k) for integer top, k >= 0.
+25 Uses falling factorial: top * (top-1) * ... * (top - k + 1) / k!.
+26 Equals math.comb when top >= 0, and uses the extension when top < 0.
+27 """
+28 if k < 0:
+29 return 0
+30 if top >= 0:
+31 if top < k:
+32 return 0
+33 return comb(top, k)
+34 # top < 0: use falling factorial. C(top, k) = (-1)^k * C(-top + k - 1, k)
+35 return (-1) ** k * comb(-top + k - 1, k)
+36
+37
+38def bala_rhs(n: int, m: int) -> Fraction:
+39 """Compute Peter Bala's conjectured formula RHS:
+[…]
+41 """
+42 s = Fraction(0)
+43 for k in range(1, n + 1):
+44 sign = 1 if k % 2 == 1 else -1
+45 term = Fraction(sign, k) * gen_binom(m * k, k) * gen_binom(n + (m - 1) * k, n - k)
+46 s += term
+47 return s / m
+48
+49
+50def main() -> None:
+[…]
+52 print()
+53
+54 # m must be nonzero integer. Negative m uses generalized (falling-factorial) binomial.
+55 test_ms = [-7, -5, -4, -3, -2, -1, 1, 2, 3, 4, 5, 6, 7, 8, 10, 15]
+56 test_ns = list(range(1, 41))
+57
+58 all_ok = True
+59 for m in test_ms:
+60 for n in test_ns:
+61 lhs = harmonic(n)
+62 rhs = bala_rhs(n, m)
+63 ok = lhs == rhs
+64 if not ok:
+65 print(f" m={m:+d} n={n:2d} H(n)={lhs} RHS={rhs} [FAIL]")
+66 all_ok = False
+67 print(f" m={m:+d}: all n in 1..{test_ns[-1]} verified OK")
+68 print()
+69 print("All tests passed." if all_ok else "FAILURES detected.")
+70
+71
+72if __name__ == "__main__":
+73 main()
+Output (abridged):
+[…]
+                                          * C(mk,k) * C(n+(m-1)k, n-k)
+  m=-7: all n in 1..40 verified  OK
+  m=-5: all n in 1..40 verified  OK
+  m=-4: all n in 1..40 verified  OK
+  m=-3: all n in 1..40 verified  OK
+  m=-2: all n in 1..40 verified  OK
+  m=-1: all n in 1..40 verified  OK
+  m=+1: all n in 1..40 verified  OK
+  m=+2: all n in 1..40 verified  OK
+  m=+3: all n in 1..40 verified  OK
+  m=+4: all n in 1..40 verified  OK
+  m=+5: all n in 1..40 verified  OK
+  m=+6: all n in 1..40 verified  OK
+  m=+7: all n in 1..40 verified  OK
+  m=+8: all n in 1..40 verified  OK
+  m=+10: all n in 1..40 verified  OK
+  m=+15: all n in 1..40 verified  OK
+All tests passed.
+The script depends only on the Python standard library (fractions.Fraction and math.comb).
+The total runtime is a fraction of a second on commodity hardware.
+Generalised binomial coefficients $\binom{a}{b}$ for negative $a$ are evaluated by the falling-factorial extension $\binom{-k}{j}=(-1)^{j}\binom{k+j-1}{j}$, matching the convention of Section 1.
+Appendix B Formal-power-series verification of the four intermediate identities
+This second script verifies, by truncating to degree $N=8$ and comparing coefficients, the four formal-power-series identities used in the proof of Theorem 1:
+  1. (i)
+    the Lagrange identity (5) of Lemma 2,
+  2. (ii)
+    its sign-alternating version (7),
+  3. (iii)
+    the classical evaluation $\sum_{k\geq 0}\binom{mk}{k}u^{k}=(1+w)/(1+(1-m)w)$ of (9), and
+  4. (iv)
+    the substitution $V(x)=x$ of Lemma 4.
+For each $m\in\{-3,-2,-1,1,2,3,4,5\}$, the implicit fixed-point equation $w=u(1+w)^{m}$ is solved iteratively as a truncated power series, and the four identities are then verified coefficient-by-coefficient.
+All $4\times 8=32$ instances pass.
+Listing 2: check_proof.py: formal-power-series verification of the four intermediate identities.
+1"""
+2Sanity-check the key identities used in the proof of Bala's harmonic-number conjecture:
+3
+[…]
+[…]
+6 where v = u (1-v)^m.
+7
+8Both checked as power-series identities, by truncating to degree N and comparing
+9coefficients of [u^k].
+10
+11We also verify the key substitution: if u = x/(1-x)^m, then the power-series
+12solution v(u) of v = u(1-v)^m satisfies v = x, and so
+13
+[…]
+15
+16confirming Bala's identity S_m(n) = H_n.
+17"""
+18
+19from fractions import Fraction
+20from math import comb
+21
+22
+23def gen_binom(top: int, k: int) -> int:
+24 """Generalized binomial C(top, k) for integer top and k >= 0.
+25 Uses falling factorial extension for negative top."""
+26 if k < 0:
+27 return 0
+28 if top >= 0:
+29 if top < k:
+30 return 0
+31 return comb(top, k)
+32 return (-1) ** k * comb(-top + k - 1, k)
+33
+34
+35def truncate_log1p(series_coeffs: list[Fraction], N: int) -> list[Fraction]:
+36 """Given coefficients of f(u) = sum c_k u^k with c_0 = 0,
+37 compute coefficients of log(1 + f(u)) up to degree N.
+38 """
+39 # Use formal power series logarithm via integration of f'/(1+f).
+40 # Easier: log(1+f) where f = sum c_k u^k, c_0 = 0.
+41 # Method: g = log(1+f) satisfies (1+f) g' = f' so we can extract coefs recursively:
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+49 f = list(series_coeffs)
+50 while len(f) <= N:
+51 f.append(Fraction(0))
+52 d = [Fraction(0)] * (N + 1)
+53 for n in range(1, N + 1):
+54 s = Fraction(0)
+55 for i in range(1, n):
+56 s += f[i] * (n - i) * d[n - i]
+57 d[n] = f[n] - s / n
+58 return d
+59
+60
+61def lagrange_solve(phi: list[Fraction], N: int) -> list[Fraction]:
+62 """Given phi = [phi_0, phi_1, ...] coefficients of a power series with phi_0 != 0,
+[…]
+64 truncated to degree N.
+65
+66 Uses the recursion implied by w = u * phi(w): write w = u * phi(w) and
+67 repeatedly substitute. We use the formal power series approach.
+68 """
+[…]
+70 w = [Fraction(0)] * (N + 1)
+71 for _ in range(N + 1):
+72 # phi(w) up to degree N (composition).
+73 phi_w = [Fraction(0)] * (N + 1)
+74 # phi_w = sum_j phi_j * w^j
+75 wpow = [Fraction(0)] * (N + 1)
+76 wpow[0] = Fraction(1) # w^0 = 1
+77 for j in range(len(phi)):
+78 for d in range(N + 1):
+79 phi_w[d] += phi[j] * wpow[d]
+80 # Update wpow = wpow * w (truncated)
+81 new_wpow = [Fraction(0)] * (N + 1)
+82 for a in range(N + 1):
+83 if wpow[a] == 0:
+84 continue
+85 for b in range(N + 1 - a):
+86 new_wpow[a + b] += wpow[a] * w[b]
+87 wpow = new_wpow
+88 # New w = u * phi_w, i.e., shift right by 1.
+89 new_w = [Fraction(0)] * (N + 1)
+90 for d in range(N):
+91 new_w[d + 1] = phi_w[d]
+92 w = new_w
+93 return w
+94
+95
+96def series_mul(a: list[Fraction], b: list[Fraction], N: int) -> list[Fraction]:
+97 out = [Fraction(0)] * (N + 1)
+98 for i in range(min(len(a), N + 1)):
+99 if a[i] == 0:
+100 continue
+101 for j in range(min(len(b), N + 1 - i)):
+102 out[i + j] += a[i] * b[j]
+103 return out
+104
+105
+106def series_inverse(a: list[Fraction], N: int) -> list[Fraction]:
+107 """Multiplicative inverse of a power series with a[0] != 0."""
+108 inv = [Fraction(0)] * (N + 1)
+109 inv[0] = 1 / a[0]
+110 for n in range(1, N + 1):
+111 s = Fraction(0)
+112 for i in range(1, min(n, len(a) - 1) + 1):
+113 s += a[i] * inv[n - i]
+114 inv[n] = -s / a[0]
+115 return inv
+116
+117
+118def main() -> None:
+119 N = 8
+120 print(f"Verifying identities to degree {N}.")
+121 print()
+122
+123 for m in (-3, -2, -1, 1, 2, 3, 4, 5):
+[…]
+125 lhs = [Fraction(0)] * (N + 1)
+126 for k in range(1, N + 1):
+127 lhs[k] = Fraction(1, k) * gen_binom(m * k, k)
+128
+129 # Compute w with phi(w) = (1+w)^m (coefficients), truncated to deg N.
+130 # For m >= 0, phi is polynomial of degree m. For m < 0, it's an infinite series.
+131 phi_len = (m + 1) if m >= 0 else (N + 1)
+132 phi = [Fraction(gen_binom(m, j)) for j in range(phi_len)]
+133 w = lagrange_solve(phi, N)
+134
+135 # RHS: m * log(1 + w).
+136 log1pw = truncate_log1p(w, N)
+137 rhs = [m * c for c in log1pw]
+138
+139 ok = lhs == rhs
+140 print(f" m={m} Identity (I) [sum (1/k) C(mk,k) u^k = m log(1+w)] OK={ok}")
+141 if not ok:
+142 print(f" LHS: {lhs}")
+143 print(f" RHS: {rhs}")
+144
+145 # Check identity (II): sum (-1)^(k+1)/k * C(mk,k) u^k = -m log(1-v) with v = u(1-v)^m.
+146 lhs2 = [Fraction(0)] * (N + 1)
+147 for k in range(1, N + 1):
+148 lhs2[k] = Fraction((-1) ** (k + 1), k) * gen_binom(m * k, k)
+149 # phi2(v) = (1-v)^m, truncated to deg N for m < 0.
+150 phi2_len = (m + 1) if m >= 0 else (N + 1)
+151 phi2 = [Fraction((-1) ** j * gen_binom(m, j)) for j in range(phi2_len)]
+152 v = lagrange_solve(phi2, N)
+153 # log(1 - v) = log(1 + (-v)).
+154 neg_v = [-c for c in v]
+155 log1mv = truncate_log1p(neg_v, N)
+156 rhs2 = [-m * c for c in log1mv]
+157 ok2 = lhs2 == rhs2
+158 print(f" m={m} Identity (II) [sum (-1)^(k+1)/k C(mk,k) u^k = -m log(1-v)] OK={ok2}")
+159 if not ok2:
+160 print(f" LHS: {lhs2}")
+161 print(f" RHS: {rhs2}")
+162
+163 # Check that when u = x/(1-x)^m, v = x.
+164 # This is a substitution check at the formal level: use power series in x.
+165 Nx = N
+166 u_of_x = [Fraction(0)] * (Nx + 1)
+[…]
+168 for j in range(Nx):
+169 u_of_x[j + 1] = Fraction(gen_binom(m + j - 1, j))
+170 # Compute v_of_x via composition v(u(x)).
+171 # We need v as a power series in u, then substitute u = u_of_x.
+172 # v(u) = sum v_k u^k where v from above.
+173 # Compose: v_of_x = sum v_k * (u_of_x)^k.
+174 v_of_x = [Fraction(0)] * (Nx + 1)
+175 upowx = [Fraction(0)] * (Nx + 1)
+176 upowx[0] = Fraction(1)
+177 for k in range(Nx + 1):
+178 for d in range(Nx + 1):
+179 v_of_x[d] += v[k] * upowx[d]
+180 new_upowx = [Fraction(0)] * (Nx + 1)
+181 for a in range(Nx + 1):
+182 if upowx[a] == 0:
+183 continue
+184 for b in range(Nx + 1 - a):
+185 new_upowx[a + b] += upowx[a] * u_of_x[b]
+186 upowx = new_upowx
+187 target = [Fraction(0)] * (Nx + 1)
+188 target[1] = Fraction(1) # v = x
+189 ok3 = v_of_x == target
+190 print(f" m={m} Substitution u = x/(1-x)^m gives v = x OK={ok3}")
+191 if not ok3:
+192 print(f" v_of_x: {v_of_x}")
+193
+[…]
+195 bcoef = [Fraction(gen_binom(m * k, k)) for k in range(N + 1)]
+196 one_plus_w = [Fraction(0)] * (N + 1)
+197 one_plus_w[0] = Fraction(1)
+198 for k in range(N + 1):
+199 one_plus_w[k] += w[k]
+200 one_plus_1mw = [Fraction(0)] * (N + 1)
+201 one_plus_1mw[0] = Fraction(1)
+202 for k in range(N + 1):
+203 one_plus_1mw[k] += (1 - m) * w[k]
+204 denom_inv = series_inverse(one_plus_1mw, N)
+205 rhs5 = series_mul(one_plus_w, denom_inv, N)
+206 ok5 = bcoef == rhs5
+207 print(f" m={m} Identity (5) [sum C(mk,k) u^k = (1+w)/(1+(1-m)w)] OK={ok5}")
+208 if not ok5:
+209 print(f" LHS: {bcoef}")
+210 print(f" RHS: {rhs5}")
+211 print()
+212
+213
+214if __name__ == "__main__":
+215 main()
+Output (abridged):
+Verifying identities to degree 8.
+  m=-3  Identity (I)  [sum (1/k) C(mk,k) u^k = m log(1+w)]    OK=True
+  m=-3  Identity (II) [sum (-1)^(k+1)/k ...     = -m log(1-v)] OK=True
+  m=-3  Substitution u = x/(1-x)^m gives v = x                 OK=True
+  m=-3  Identity (5)  [sum C(mk,k) u^k = (1+w)/(1+(1-m)w)]     OK=True
+  m=+5  Identity (I)                                           OK=True
+  m=+5  Identity (II)                                          OK=True
+  m=+5  Substitution u = x/(1-x)^m gives v = x                 OK=True
+  m=+5  Identity (5)                                           OK=True
+The script depends only on the Python standard library.
+Power-series multiplication, multiplicative inverse, formal logarithm, and the Lagrange-type fixed-point solver ($w=u\,\phi(w)$ for $\phi$ a polynomial or formal series) are all implemented from scratch in $\sim 100$ lines, so the verification itself has no third-party dependencies and is straightforward to audit.